4. (9x − 4y + 4)dx − (2x − y +) = 0


1
Expert's answer
2021-05-10T13:03:46-0400

Given equation is-

"(9x-4y+4)dx-(2x-y+1)dy=0"


Let "x=u\Rightarrow dx=du"

"y=v+1\Rightarrow dy=dv"


"(9u-4v)du-(2u-v)dv=0"


Let "u=vz \Rightarrow du=vdz"

"\Rightarrow (9vz-4v)(zdv+vdz)-(2vz-v)dv=0\\[9pt]



\Rightarrow (9z-4)(z dv+vdz)-(2z-1)dz=0

\\[9pt]

\Rightarrow(9z^2-6z+1)dv+v(9z-4)dz=0

\\[9pt]

\Rightarrow \dfrac{dv}{v}+\dfrac{9z-4}{(3z-1)^2}dz=0

\\[9pt]

\Rightarrow \dfrac{dv}{v}+\dfrac{3dz}{3z-1}-\dfrac{dz}{(3z-1)^2}=0"


Integrating Both the sides-


"lnv+ln|3z-1|+\dfrac{1}{9z-3}+lnc=0\\[9pt]

\Rightarrow 1=-3(3z-1)ln|c(3zv-v)|"




"v=-3(3vz-v)ln|c(3zv-v)|"


Putting "z=\dfrac{u}{v} , u=x \text{ and }v=y-1"


so we get-

"y=1+3(y-3x-1)ln|c(3x-y+1)|"